# A Formal Comparison Between Chain-of-Thought and Latent Thought

### v1

@@ C.3.3 Feedforward layers
where ${\bm{e}}_{i}\in\{0,1\}^{n}$. By Lemma C.6, when substituting ${\bm{x}}=(\bm{e}(x_{1}),\ldots,\bm{e}(x_{n}))$ into the lemma, there exists a feed-forward layer $\mathrm{FF}_{1}$ such that
@@
Then, we show that the feed-forward layer can encode the entire computation graph into its weights and simulate all nodes simultaneously. Let the flag vector for each node be $(t_{1},\ldots,t_{N})\in\{0,1\}^{N}$, where $N\coloneqq|V_{n}|$. By Lemmas C.7 and C.8, there exist feed-forward layers $\mathrm{FF}_{2},\mathrm{FF}_{3}:\mathbb{F}_{s}^{N(|\Sigma|+1)}\to\mathbb{F}_{s}^{N(|\Sigma|+1)}$ such that, for the input vector $(z_{1},\ldots,z_{N})\in\Sigma^{N}.$,
@@
To extract the output node corresponding to each position denoted by $o_{i}$, in the final loop iteration, by Lemma C.6, there exists a feedforward layer $\mathrm{FF}_{4}$ such that
@@
We construct the looped model as follows. Each input token $x_{i}\in\Sigma$ at position $i\in[n]$ is embedded as ${\bm{h}}^{(0)}_{i}\in\{0,1\}^{2n+(1+|\Sigma|)N+|\Sigma|}$:
${\bm{h}}^{(0)}_{i}=\bigl(\,{\bm{e}}_{i},\;{\bm{e}}_{o_{i}},\;\bm{e}(x_{i}),\;0,\;\bm{e}(x_{i}),\;0,\;\ldots,\;\bm{e}(x_{i}),\;0,\;{\bm{0}}_{(1+|\Sigma|)(N-n)+|\Sigma|}\bigr).$ (53)
@@
The parameter size grows proportionally with the input dimension $\mathrm{size}(G_{n})$, since the function must be approximated along each dimension. Therefore, it can be bounded by $O\!\left(\mathrm{ff\_param}(G_{n})\cdot\mathrm{size}(G_{n})\right).$Discussion: One might object that our proof, compressing the entire length $n$ input to a single position and then applying an arbitrary feed-forward computation, deviates from standard Transformer architecture. An alternative is to distribute computation across positions, as shown by Sanford et al. (2024b), which can reduce the embedding dimension. We nevertheless adopt the single-position construction to isolate the essence of looped models: attention serves solely to aggregate information, a single layer is sufficient, while the feed-forward network performs the computation in latent space, which is more expressive than computations in the language space. This perspective aligns with expressivity results for looped ReLU networks (Liang et al., 2024).
C.6 Proof for Theorem 3.13
@@
For the upper bound $\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]\subseteq\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k}),$ we follow the argument of Li et al. (2024). Their key observation is that a restricted form of automaton can model iterative computation under constant precision: the rounding operation preserves monotonicity, and constant precision yields counter-free, restricted state spaces, which are therefore computable by $\mathsf{AC}^{0}$ circuits. In the case of polynomial precision, prefix summation can be simulated by $\mathsf{TC}^{0}$, which also allows the detection and correction of rounding in fixed-point arithmetic.
Theorem C.9 (Li et al., 2024).
@@
$\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]\subseteq\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k}):$ In Transformers, constant-depth computation is defined by Summation with Iterative Rounding (see Definition C.3), which by Theorem C.9 can be simulated in $\mathsf{AC}^{0}$ (resp. $\mathsf{TC}^{0}$). A Looped TF simply stacks these computations vertically through iteration, and thus the result follows.
@@
$\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k})\subseteq\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]:$ Since Boolean circuits are DAGs, the claim follows directly from Theorem 3.11 together with Lemma C.10. ∎
Appendix D Deferred Proofs for Section 4
D.1 Preliminaries
In this section, we state the deferred definitions and present the lemma.
D.1.1 Models of Computation
To analyze the distinction between CoT and latent thought in the context of next-token prediction, we formalize CoT as a process that, given the prefix $y_{<i}=(y_{1},y_{2},\ldots,y_{i-1})$, generates the next token $y_{i}$ via a stochastic reasoning step, as follows.
Definition D.1 (CoT with stochastic decoding).
@@ D.1.1 Models of Computation
Let $\mathcal{V}$} be a vocabulary. Given an input $x\in{\mathcal{V}}^{*}$, CoT stochastically generates a sequence of output blocks of the form
$\Big(r_{1},\,e,\,y_{1},\,e^{\prime},\;r_{2},\,e,\,y_{2},\,e^{\prime},\;\cdots\;r_{m},\,e,\,y_{m},\,e^{\prime}\Big)$
@@
where each $r_{i}\in{\mathcal{V}}^{*}$ is a reasoning path and $e,e^{\prime}\in{\mathcal{V}}$ are special delimiter tokens, yielding the final output string $y_{1}\dots y_{m}$. Generation proceeds autoregressively via next-token prediction: for each $i\in[m]$, the model generates a reasoning step $r_{i}$ followed by an output token $y_{i}$, conditioned on the input $x$, previous outputs $y_{<i}$, and prior reasoning steps $r_{<i}$.
For looped models, we use the last output of Looped TF as the next token.
Definition D.2 (Looped TF with stochastic decoding).
Let $f$ be a Looped TF with vocabulary $\Sigma$ and loop iteration bound $T(n)\in\mathsf{poly}(n)$. Given an input $x\in\Sigma^{*}$ and prefix $y_{<k}=(y_{1},\dots,y_{k-1})$, its hidden representation at the last position after $T(|x|)$ iterations is
$h(x,y_{\leq k-1})\coloneq f^{T(|x|)}(x,y_{\leq k-1})_{n+k-1},$
A stochastic decoding function $\mathrm{Dec}:\mathbb{F}^{d}\to\Delta(\Sigma)$ maps the hidden state at each position to a probability distribution over tokens. The autoregressive generation distribution of the model is then
$\Pr(y\mid x)\;\coloneq\;\prod_{k=1}^{m}\mathrm{Dec}\bigl(h(x,y_{<k})\bigr)(y_{k}),$
for every $y=(y_{1},\dots,y_{m})\in\Sigma^{*}$.
D.1.2 Self-Consistency Enhances Reliability
The lemma shows that the success probability can be improved by repeating the trial multiple times.
Lemma D.3 (Error reduction).
Suppose there is a randomized estimator $A$ for a nonnegative function $f$ such that for every input $x$, every accuracy parameter $\varepsilon>0$, and some constant $\gamma\in(0,1/2)$,
$\Pr\!\left[(1-\varepsilon)f(x)\ \leq\ A(x,\varepsilon)\ \leq\ (1+\varepsilon)f(x)\right]\ \geq\ \tfrac{1}{2}+\gamma,$ (63)
and the running time of $A$ is polynomial in $|x|$ and $1/\varepsilon$. Then, for every confidence parameter $\delta\in(0,1)$, there is an FPRAS for $f$ outputs $\widehat{f}(x)$ satisfying
$\Pr\!\left[(1-\varepsilon)f(x)\ \leq\ \widehat{f}(x)\ \leq\ (1+\varepsilon)f(x)\right]\ \geq\ 1-\delta.$ (64)

### v2

@@ B.3.3 Feedforward layers
where ${\bm{e}}_{i}\in\{0,1\}^{n}$. By Lemma B.6, when substituting ${\bm{x}}=(\bm{e}(x_{1}),\ldots,\bm{e}(x_{n}))$ into the lemma, there exists a feed-forward layer $\mathrm{FF}_{1}$ such that
@@
Then, we show that the feed-forward layer can encode the entire computation graph into its weights and simulate all nodes simultaneously. Let the flag vector for each node be $(t_{1},\ldots,t_{N})\in\{0,1\}^{N}$, where $N\coloneqq|V_{n}|=\mathrm{size}(G_{n})$. By Lemmas B.7 and B.8, there exist feed-forward layers $\mathrm{FF}_{2},\mathrm{FF}_{3}:\mathbb{F}_{s}^{N(|\Sigma|+1)}\to\mathbb{F}_{s}^{N(|\Sigma|+1)}$ such that, for the input vector $(z_{1},\ldots,z_{N})\in\Sigma^{N}.$,
@@
To extract the output node corresponding to each position denoted by $o_{i}$, in the final loop iteration, by Lemma B.6, there exists a feedforward layer $\mathrm{FF}_{4}$ such that
@@
We construct the looped model as follows. Each input token $x_{i}\in\Sigma$ at position $i\in[n]$ is embedded as
${\bm{h}}^{(0)}_{i}=\bigl(\,{\bm{e}}_{i},\;{\bm{e}}_{o_{i}},\;\bm{e}(x_{i}),\;0,\;\bm{e}(x_{i}),\;0,\;\ldots,\;\bm{e}(x_{i}),\;0,\;{\bm{0}}_{(1+|\Sigma|)(N-n)+|\Sigma|}\bigr)\in\{0,1\}^{2n+(1+|\Sigma|)N+|\Sigma|}.$ (65)
@@
The parameter size grows proportionally with the input dimension $\mathrm{size}(G_{n})$, since the function must be approximated along each dimension. Therefore, it can be bounded by $O\!\left(\mathrm{ff\_param}(G_{n})\cdot\mathrm{size}(G_{n})\right).$
@@
Discussion: Our proof compresses the entire input into a single position before applying an arbitrary feed-forward computation, which may appear to deviate from the standard Transformer architecture. An alternative approach is to distribute computation across multiple positions, as shown by (Sanford et al., 2024b), which can reduce the required embedding dimension. We nevertheless adopt the single-position construction to isolate the core characteristics of looped models: attention is used purely as an information aggregation mechanism, a single Transformer layer suffices, and all computation is carried out by the feed-forward network in latent space. This latent-space computation is strictly more expressive than computation in the language space. This is supported by recent work for looped ReLUs (Liang et al., 2024).
B.6 Proof of Theorem 3.6 for Continuous Thought
Proof.
The proofs are based on the construction for CoT and looped TF. Let the vocabulary be ${\mathcal{V}}=\Sigma$ and each node $v\in V_{n}$ is labeled by a one-hot vector $\bm{e}(v)\in\{0,1\}^{|\mathcal{F}|}$. At decoding step $k$, the model has access to the concatenated sequence
$(x_{1},x_{2},\ldots,x_{n},\,h_{1},h_{2},\ldots,h_{k})$ (75)
where $x=(x_{1},\ldots,x_{n})\in\Sigma^{n}$ denotes the input, and $h_{i}\in\mathbb{F}_{s}^{m}$ is the hidden state generated at the $i$-th Coconut step. For each node $v_{j}$, let $v_{j}(x)$ denote its value on input $x$. We assume that
$h_{k}\coloneq\bigl(t_{1,k-1}\cdot\bm{e}(v_{1}(x)),\,t_{1,k},\,t_{2,k-1}\cdot\bm{e}(v_{2}(x)),\,t_{2,k},\,\ldots,\,t_{N,k-1}\cdot\bm{e}(v_{N}(x)),\,t_{N,k},\,{\bm{e}}^{\prime}_{k},\,{\bm{0}}_{|\Sigma|}\bigr),$
where $N\coloneqq\mathrm{size}(G_{n})$, $t_{i,k}\in\{0,1\}$ indicates whether node $v_{i}$ lies within depth at most $k$, and $\bm{e}:\Sigma\to\{0,1\}^{|\Sigma|}$ denotes one-hot encoding, and the vector ${\bm{e}}^{\prime}_{k}\in\{0,1\}^{N}$ is defined by
${\bm{e}}^{\prime}_{k}=\begin{cases}\mathbf{0}&k<\mathrm{depth}(G_{n}),\\[4.0pt] \bm{e}(v_{o_{i}})&k=\mathrm{depth}(G_{n})+i,\end{cases}$ (76)
where $v_{o_{i}}$ denotes the $i$-th output node, and ${\bm{e}}^{\prime}_{k}$ can be encoded by positional embeddings. Under this assumption, we prove by induction that the model generates $h_{k+1}(x)$.
We first consider the base case $k=1$, in which the model receives only $\bm{x}$}. We focus on the final token. Using the same construction as in the looped TF, we can show that a single feed-forward layer followed by an attention layer suffices to copy all input tokens to every position. Specifically, the hidden representation at the last position becomes
$h_{1}=\bigl(\bm{e}(v_{1}(x)),\,1,\,\ldots,\,\bm{e}(v_{n}(x)),\,1,\,{\bm{0}},\ 1,\,{\bm{0}},\,{\bm{e}}^{\prime}_{0},\,{\bm{0}}\bigl),$ (77)
where $v_{i}(x)=x_{i}$ for all $i\in[n]$, corresponding to the input nodes. In what follows, we consider the case $k>1$. By the same argument as for Looped TF, based on Equation 62, we show that the feed-forward layers can encode the entire computation graph into their weights and simulate all nodes simultaneously. That is, there exist two feed-forward layers whose composition, denoted by $\mathcal{FF}$, satisfies $\mathcal{FF}(h_{k})=h_{k+1}.$
Consider the last feed-forward layer applied after the zero-weight attention layer. By substituting ${\bm{e}}_{i}={\bm{e}}_{k}^{\prime}$ and ${\bm{x}}={(h_{k})}_{1:m-|\Sigma|}$ into Lemma B.6, there exists a feed-forward network $\mathrm{FF}$ such that
$(\operatorname{id}+\mathrm{FF})(h_{k})=\begin{cases}h_{k+1},&\text{if }k<\mathrm{depth}(G_{n}),\\[6.0pt] \Bigl({(h_{k+1})}_{1:m-|\Sigma|},\,\bm{e}(v_{o_{i}}(x))\Bigr),&\text{if }k=\mathrm{depth}(G_{n})+i.\end{cases}$ (78)
In particular, for all $k<\mathrm{depth}(G_{n})$, the output of the last token, for the input $(x_{1},x_{2},\ldots,x_{n},\;h_{1},h_{2},\ldots,h_{k})$ is $h_{k+1}$. For the final positions corresponding to output nodes, namely $k=\mathrm{depth}(G_{n})+i$, the hidden state ${\bm{h}}_{k}$ contains the embedding of the output symbol $v_{o_{i}}(x)$. Applying the output projection yields
$\mathbf{OUT}({\bm{h}}_{k})=\begin{bmatrix}\mathbf{0}&{\bm{I}}_{|\Sigma|}\end{bmatrix}{\bm{h}}_{k}=\bm{e}\!\left(v_{o_{i}}(x)\right).$ (79)
Finally, the decoding function outputs the symbol in $\Sigma$ corresponding to the maximum coordinate of $\mathbf{OUT}({\bm{h}}_{k})$. ∎
B.7 Proof for Theorem 3.12
@@
For the upper bound $\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]\subseteq\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k}),$ we follow the argument of (Li et al., 2024). Their key observation is that a restricted form of automaton can model iterative computation under constant precision: the rounding operation preserves monotonicity, and constant precision yields counter-free, restricted state spaces, which are therefore computable by $\mathsf{AC}^{0}$ circuits. In the case of polynomial precision, prefix summation can be simulated by $\mathsf{TC}^{0}$, which also allows the detection and correction of rounding in floating-point arithmetic.
Theorem B.9 (Li et al., 2024).
@@
$\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]\subseteq\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k}):$ In Transformers, constant-depth computation is defined by Summation with Iterative Rounding (see Definition B.3), which by Theorem B.9 can be simulated in $\mathsf{AC}^{0}$ (resp. $\mathsf{TC}^{0}$). A looped TF simply stacks these computations vertically through iteration, and thus the result follows.
@@
$\mathsf{AC}^{k}\ \textup{(resp.\ }\mathsf{TC}^{k})\subseteq\mathsf{Loop}[\log^{k}n,\,\mathsf{poly}(n),\,1\ \textup{(resp.\ }\log n)]:$ Since Boolean circuits are DAGs, the claim follows directly from Theorem 3.10 together with Lemma B.10. ∎
Appendix C Deferred Proofs for Section 4
C.1 Definition for Models of Computation
We model a language model as a probabilistic process that, given an input and a generated prefix, produces either an internal reasoning state or an output token. This process induces a probability distribution over final outputs. We first formalize CoT under this setting. We focus on the saturated hardmax attention as in (Merrill et al., 2022; Nowak et al., 2024).
Definition C.1 (Language model with CoT).
@@ C.1 Definition for Models of Computation
Let $\mathcal{V}$} be a vocabulary. Given an input $x\in{\mathcal{V}}^{*}$, a language model with CoT stochastically generates a sequence of output blocks of the form
$\Big(r_{1},\,e,\,y_{1},\,e^{\prime},\;r_{2},\,e,\,y_{2},\,e^{\prime},\;\cdots\;r_{m},\,e,\,y_{m},\,e^{\prime}\Big),$ (80)
@@
where each $r_{i}\in{\mathcal{V}}^{*}$ represents an explicit reasoning trace, $y_{i}\in{\mathcal{V}}$ is an output token, and $e,e^{\prime}\in{\mathcal{V}}$ are special delimiter tokens. The final output is the string $y_{1}\cdots y_{m}$. Generation proceeds autoregressively: at iteration $i$, the model generates a reasoning segment $r_{i}$ followed by an output token $y_{i}$, conditioned on the input $x$, the previously generated outputs $y_{<i}$, and prior reasoning segments $r_{<i}$. We denote by $p(y\mid x)$ the resulting distribution over final outputs.
Then we define the language models with latent thought reasoning: Coconut and looped TF.
Definition C.2 (Language model with Coconut).
Let $\mathcal{V}$} be a vocabulary. Given an input $x\in{\mathcal{V}}^{*}$, a language model with Coconut generates a sequence of blocks
$\Big(r_{1},\,y_{1},\,r_{2},\,y_{2},\,\ldots,\,r_{m},\,y_{m}\Big),$ (81)
where each $r_{i}\in{\mathbb{F}^{d}}^{*}$ represents an internal continuous reasoning state, and each $y_{i}\in{\mathcal{V}}$ is an output token.
Generation proceeds autoregressively. At each iteration, the internal continuous reasoning state $r_{i}$ is generated deterministically as a function of the input $x$, the previously generated outputs $y_{<i}$, and the prior internal states $r_{<i}$, while the output token $y_{i}$ is generated stochastically. The decision of when to emit an output token is determined internally by the model.
For looped TF models, the computation proceeds through internally repeated iterations before producing any output tokens. The number of iterations is determined by the model as a function of the input.
Definition C.3 (Language model with looped TF).
Given an input $x\in{\mathcal{V}}^{*}$, a looped TF produces an output sequence autoregressively. At each iteration $i\in[m]$, the model internally performs a number of repeated transformation steps before emitting an output token $y_{i}$, conditioned on the input $x$ and the previously generated outputs $y_{<i}$. The number of internal loop iterations required to generate each output token is determined internally by the model as a function of the input $x$ and the generated prefix $y_{<i}$.
We define complexity classes corresponding to language models under different reasoning paradigms. In contrast to the non-uniform models typically used in parallel computation analysis, we adopt a uniform setting analogous to Turing machines: a single model with a fixed set of parameters is applied to all input lengths, while the number of reasoning steps is allowed to grow as a function of the input size $n$. Following the convention in (Merrill and Sabharwal, 2024), we allow $O(\log n)$ bits of numerical precision to represent positional embeddings and internal activations. Furthermore, we extend the output space beyond simple binary decisions. Specifically, the model’s output is not restricted to $\Sigma^{k}$, but can be an arbitrary finite binary string in $\Sigma^{*}$. This extension enables the model to represent functions of the form $f:\Sigma^{*}\to\mathbb{R}$, thereby capturing counting, probabilistic modeling, and approximation tasks.
Definition C.4 (Complexity Classes $\mathsf{pCoT}$, $\mathsf{pCT}$, and $\mathsf{pLOOP}$).
Let $\mathsf{pCoT}[T(n)]$, $\mathsf{pCT}[T(n)]$, and $\mathsf{pLOOP}[T(n)]$ denote the classes of probabilistic computations that define an output distribution $p:\Sigma^{*}\times\Sigma^{*}\to[0,1]$. A function $f$ belongs to such a class if there exists a language model $\mathcal{M}$, employing CoT, Coconut, or looped TF, respectively, such that for any input $x\in\Sigma^{n}$, the model induces the distribution $p(x,\cdot)$ within $O(T(n))$ steps using $O(\log n)$ bits of numerical precision.
C.2 Lemma: Universality of Chain of Thought
Definition C.5.
A probabilistic Turing machine $M=(Q,\Sigma,\Gamma,q_{0},q_{1},\delta_{1},\delta_{2})$ is defined as:
  • $Q$ is a finite set of states,
  • $\Sigma$ is the input/output alphabet,
Prior work has shown that probabilistic CoT can simulate any such PTM, thereby demonstrating its universality.
Lemma C.6 (Nowak et al., 2024).
$\Pr\bigl[\mathsf{CoT}_{M}(x)=y\bigr]\;=\;\Pr\bigl[M(x)=y\bigr].$ (82)
Moreover, the number of reasoning steps of $\mathsf{CoT}_{M}$ is bounded by $\mathsf{poly}(|x|)$.
C.3 Self-Reducibility and Complexity of Approximate Counting
Here, we provide the definitions of relation and associated counting problems.
Definition C.7 (Relation).
A relation over an alphabet $\Sigma$ is a subset $R\subseteq\Sigma^{*}\times\Sigma^{*}.$ For an input $x\in\Sigma^{*}$, we denote
$R(x):=\{\,y\in\Sigma^{*}:(x,y)\in R\,\}.$
Definition C.8 (Counting).
Given a relation $R$, the associated counting function is defined as
$N_{R}:\Sigma^{*}\to\mathbb{N},\qquad N_{R}(x):=|R(x)|.$
Definition C.9 ($p$-relation).
A relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is called a $p$-relation if
  • the membership $(x,y)\in R$ can be decided in time polynomial in $|x|$, and
  • there exists a polynomial $p$ such that for all $(x,y)\in R$, we have $|y|\leq p(|x|)$.
Definition C.10.
The extension counting function associated with a relation $R\subseteq\Sigma^{*}\times\Sigma^{*}$ is
$\operatorname{EXT}_{R}:\Sigma^{*}\times\Sigma^{*}\to\mathbb{N},\qquad\operatorname{EXT}_{R}(x,w):=\bigl|\{\,z\in\Sigma^{*}:(x,wz)\in R\,\}\bigr|.$
The complexity class $\mathsf{P}$ consists of counting problems associated with $p$-relations (Valiant, 1979). Then, we provide definitions of schemes for approximate counting. In the remainder of this paper, we say that an algorithm produces an output approximating $f(x)$ within ratio $1+\varepsilon$ if its output $\hat{f}(x)$ satisfies
$(1-\varepsilon)f(x)\leq\hat{f}(x)\leq(1+\varepsilon)f(x),$
Definition C.11 (FPTAS).
An algorithm is called a fully polynomial-time approximation scheme (FPTAS) for a function $f$ if, for any input $x$ and any $\varepsilon>0$, it produces an output approximating $f(x)$ within ratio $1+\varepsilon$, and runs in time polynomial in $|x|$ and $1/\varepsilon$.
Definition C.12 (FPRAS).
An algorithm is a fully polynomial-time randomized approximation scheme (FPRAS) for a function $f$ if, for any input $x$ and any $\varepsilon>0$ and $\delta>0$, it produces an output approximating $f(x)$ within ratio $1+\varepsilon$ with probability at least $1-\delta$, and runs in time polynomial in $|x|$, $1/\varepsilon$, and $\log(1/\delta)$.
The following proposition formalizes the relationship between approximate counting and the extension counting function.
Proposition C.13 (Jerrum et al., 1986).
Let $R$ be a $p$-relation and let $x\in\Sigma^{n}$. Let $m=p(n)$ and define $r=1+\frac{\varepsilon}{2m}$. If there exists a FPRAS that approximates $\mathrm{Ext}_{R}(x,w)$ within ratio $r$ for all prefixes $w$ with probability at least $1-\delta/m$, then there exists an FPRAS that approximates $N_{R}(x)$ within ratio $1+\varepsilon$ with probability at least $1-\delta$.
This reduction transforms the global counting problem into a sequence of local estimation steps.
Proposition C.14.
For any $0<\varepsilon\leq 1$ and any integer $m\geq 1$, it holds that
$\left(1+\frac{\varepsilon}{2m}\right)^{m}<1+\varepsilon.$ (83)